OpenStudy (anonymous):

p-32=52/7(20-6) p-32=52/7(18-6) - can anyone help me with this I keep getting it wrong!!

OpenStudy (anonymous):

p-32=(52/7)*(20-6) = p-32=(52/7)*14= p-32=104 = p=72 p-32=(52/7)*(18-6) = p-32=(52/7)*12= p-32=400/7 = p=57.14 If my answer is correct them i wish to tell you that your mistake comes by the way you are puuting the brackets